Given the estimated linear model shown​ below, complete the following computations.y=13+4x1+6x2+3x3

a. Compute

y when x1=9, x2=31, and x3=28.

b. Compute

y when x1=20, x2=34, and x3=16.

c. Compute

y when x1=17, x2=35, and x3=26.

d. Compute  y whenx1=23, x2=18, and x3=31.

a. y=______

b.y=_________

c. y=________

d.y=___________
Answer #1

The estimated linear model
y=13+4x1+6x2+3x3
a.
when x1=9, x2=31, and x3=28
then y =(13+(4*x1)+(6*x2)+(3*x3))
y =(13+(4*9)+(6*31)+(3*28))
Y =319
b.
when x1=20, x2=34, and x3=16
then y = (13+(4*x1)+(6*x2)+(3*x3))
y = (13+(4*20)+(6*34)+(3*16))
y =345

c.
when x1=17, x2=35, and x3=26.
then y =(13+(4*x1)+(6*x2)+(3*x3))
y =(13+(4*17)+(6*35)+(3*26))
y =369

d.
when x1=23, x2=18, and x3=31.
then y =(13+(4*x1)+(6*x2)+(3*x3))
y =(13+(4*23)+(6*18)+(3*31))
y =306